Find the equation of the line which passes through the point (9,5

9,5) and has slope 1

1. Give your answer in the form y=m x+b

y=m x+b. The constants should be given exactly, e.g., 1/3 rather than 0.33.


1
Expert's answer
2021-10-20T13:40:40-0400

Let "y=kx+b" - equation of the given line

k is a slope, and it is equal to 1

"y=x+b"

To find the value of, we can substitute x = 9.5, y = 9.5 since we know the line passes through this point

"9.5=9.5+b \to b = 0"

Equation of the given line is "y =x"